OpenStudy (anonymous):

solve the equation x/5+6=9

hartnn (hartnn):

u can start by subtracting 6 from both sides...

OpenStudy (anonymous):

x/5+6=9 x/5+6=9-6 x/5=3 x/5*5=3*5 x=15 Im pretty sure I did this correct

OpenStudy (anonymous):

subtract 6 from both sides:\[{x\over5}+6\color{red}-\color{red}6=9\color{red}-\color{red}6~~~~~~\rightarrow~~~~~~~~{x\over5}+0=3\] now multiply both sides by 5 to get rid of the denominator:\[\color{red}{5\over1}\times{x\over5}=3\times\color{red}{5\over1}~~~~~~~~~~\rightarrow~~~~~~\large\color{blue}{x=15}\]

OpenStudy (anonymous):

ok the 5 over 1 in that problem...where did that come from??

OpenStudy (anonymous):

Reciprocal...x=1

OpenStudy (anonymous):

you multiply by the reciprocal to get rid of the denominator...